The Pear company sells a smart phone for $250. Its sales have averaged 8,000 units per month over the last year. Recently, its closest competitor Banana company reduced the price of its smart phone from $350 to $300. As a result, Pear’s sales declined by 1,500 units per month. (a) What is the cross price elasticity of demand between the Pear and Banana smart phone? Use the averaging formula. What does this indicate about the relationship between the two products? (b) If the Pear company knows that the price elasticity of demand for its phone is -1.5, what price would the Pear company have to charge to sell the same number of units as it did before the Banana company price cut? Assume that Banana company holds its price of its phone constant at $300. Use the averaging formula.


1
Expert's answer
2022-01-10T09:56:42-0500

a)"E=\frac{1500-8000}{300-350}\times\frac{300+350}{1500+8000}=8.89"


E>0, These are substitute products


b)"-1.5=\frac{Q2-1500}{300-350}"


Q2=1575